OpenStudy (sammyalabamy):

How many grams of Mg are needed to react completely with 3.80 L of a 2.50 M HF solution? Mg + 2HF yields MgF2 + H2

OpenStudy (sammyalabamy):

@Photon336 can you help me with this last one please?

OpenStudy (photon336):

this is tricky but the information we need is hidden in the information

OpenStudy (photon336):

\[Molarity_{HF}*(Volume_{HF}) = Moles~HF\]

OpenStudy (photon336):

the molarity is the number of moles per liter. if we know the volume we can easily get the number of moles of HF by multiplying the following. \[\frac{ Moles }{ L }*L = Moles_{HF}\] So try this and tell me what you get

OpenStudy (sammyalabamy):

well the L would cancel each other out

OpenStudy (photon336):

exactly

OpenStudy (photon336):

and we would be left with the number of moles of HF but we are not done quite yet

OpenStudy (photon336):

try to see if you can fin the number of moles of HF @sammyalabamy

OpenStudy (sammyalabamy):

it would be 2.5

OpenStudy (photon336):

\[2.50(\frac{ Moles }{ L })*3.80L = 9.5~moles \]

OpenStudy (sammyalabamy):

Thank you so much @Photon336

OpenStudy (photon336):

that's just the first part @sammyalabamy now we need to find the number of moles of Mg. so now. my question is the reaction balanced and how do you know?

OpenStudy (sammyalabamy):

there are 2 of H and F on each side as well as only 1 Mg

OpenStudy (photon336):

the reaction is balanced because the same number of moles of each atom are on both sides, this is important because we can only do stoichiometry with balanced equations. now let's look at the molar ratio of HF to Mg we know that we need 2Moles of HF to react with 1 mole of Mg right?

OpenStudy (sammyalabamy):

yes

OpenStudy (photon336):

so now to find out how many moles of magnesium we need we multiply the number of moles by the molar ratio of the two. 9.5moles HF *( 1 mole Mg/ 2 moles HF = 4.8 moles of Mg needed

OpenStudy (sammyalabamy):

thanks